Question #3

Given the following information, calculate the firm’s WACC. Assume the interest on the debt is 100% tax deductible.

Tax rate:                                              20%

Debt rate:                                            4%

Preferred stock dividend rate:             8% of $100 par value

Risk-free rate of return:                      2%

Market rate of return:                          11%

Stock beta:                                          1.3

Debt value:                                          $20,000,000

P/S value:                                            $5,000,000

C/S value:                                            $25,000,000

*** what information do you need from the balance sheet and income statement in order to answer the question above? I was unable to include the financials because the system states that it is too long. Thanks,
Answer #1
WACC=(Ke*E/V)+(Kp*P/V)+(D/V)*Kd*(1-Tax rate)
Where
Ke=Cost of equity
Ke=Rf+β*(Rm-Rf)
Rf=Risk-free rate of return=2%
β=Stock beta=1.3
Rm=Market rate of return=11%
Ke=2%+1.3*(11%-2%)=2%+1.3*9%=2%+11.7%=13.7%
E=Market value of equity=C/S value=$ 25000000
V=C/S value+P/S value+Debt value=25000000+5000000+20000000=$ 50000000
Kp=Cost of P/S=Preferred stock dividend rate=8%
P=P/S value=$ 5000000
D=Debt value=$ 20000000
Kd=Cost of debt=4%
Tax rate=20%
WACC=(13.7%*25000000/50000000)+(8%*5000000/50000000)+(20000000/50000000)*4%*(1-0.20)
WACC=(13.7%*0.50)+(8%*0.10)+(0.40)*4%*(0.80)
WACC=6.85+0.80+1.28=8.93%